The two rails of a railway track, insulated from each other and from the ground, are connected to a millivoltmeter. What will be the reading of the millivoltmeter when a train travels on the track at a speed of 180 km/h? The vertical component of earth’s magnetic field is 0.2 × 10^-4 T and the rails are separated by 1 m.

Grade:upto college level

1 Answers

Jitender Pal
9 years ago
Sol. v = 180 km/h = 50 m/s B = 0.2 × 10^–4 T, L = 1 m E = Bvℓ = 0.2 I 10^–4 × 50 = 10^–3 V ∴ The voltmeter will record 1 mv.